# AfriMed-QA: A Pan-African, Multi-Specialty, Medical Question-Answering Benchmark Dataset

### v3

@@ Abstract
Tobi Olatunji1,2,4, Charles Nimo2, Abraham Owodunni1,3, Tassallah Abdullahi4,12, Emmanuel Ayodele1, Mardhiyah Sanni1,4, Chinemelu Aka1, Folafunmi Omofoye4, Foutse Yuehgoh4, Timothy Faniran4, Bonaventure F. P. Dossou4,10, Moshood Yekini4, Jonas Kemp6, Katherine Heller6, Jude Chidubem Omeke4, Chidi Asuzu MD4, Naome A. Etori4,11, Aimérou Ndiaye5, Ifeoma Okoh5, Evans Doe Ocansey5, Wendy Kinara7, Michael Best2, Irfan Essa2,6, Stephen Edward Moore8, Chris Fourie9, Mercy Nyamewaa Asiedu6♠ 1Intron, 2Georgia Institute of Technology, 3The Ohio State University, 4BioRAMP, 5Masakhane, 6Google Research, 7Kenyatta University, 8University of Cape Coast, 9SisonkeBiotik, 10MILA Quebec, 11University of Minnesota, 12Brown University tobi@intron.io
11footnotetext: Authors marked with $\spadesuit$ are senior authors.
@@ 1 Introduction
Feature AfriMed-QA BioASQ Krithara et al. (2023) MedQA-SWE Hertzberg and Lokrantz (2024) MedQA Jin et al. (2020) MedMCQA Pal et al. (2022b) PubMedQA Jin et al. (2019) MMLU Hendrycks et al. (2020) HealthSearchQA Singhal et al. (2022a)
@@
To address this gap, we introduce AfriMed-QA, 22footnotetext: Available at:
https://huggingface.co/datasets/intronhealth/afrimedqa_v2
a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias. We evaluate 30 large, small, open, closed, biomedical, and general LLMs using quantitative and qualitative approaches. While the development of the dataset is still in progress, this work establishes a foundation for acquiring diverse and representative health benchmark datasets across LMICs. The dataset is released under a CC-BY-NC-SA 4.0 license.
@@ 3.1 Data Collection
We adapted a web-based platform previously developed by Intron Health  [^1] Intron Health’s biomedical crowd-sourcing platform https://speech.intron.health to crowd-source accented and multilingual clinical speech data at scale across Africa Olatunji et al. (2023). Custom user interfaces were developed to collect each question type, for quality reviews, and for blind human evaluation of LLM responses. The MCQ User Interface and other details about the data collection tool can be found in Appendix Figure 13.
@@ 3.2 Contributor Recruitment and Instructions for Data Collection and Evaluation
Human Evaluations: Consumers provided ratings for LLM responses to CQs on relevance, helpfulness, and local context, but NOT correctness. Due of the higher expertise required, only confirmed clinicians were granted access to projects rating MCQ, SAQ, and CQ answers. Clinician status was confirmed through submitted credentials and background checks reviewing publicly available information about them. Raters were randomly assigned (double-blind) to the answer source (human or LLM). More details on human evaluations are provided in Section 4.3
@@ 4.1 LLM Selection
We evaluate 30 LLMs (Table 3) including open-source and proprietary LLMs, general-purpose and biomedical LLMs, mixture-of-experts, and models of varying sizes from 3B to over 540B parameters Chowdhery et al. (2023)). Of these, 13 are proprietary while 17 are open-source. We evaluate LLMs like Phi-3 Abdin et al. (2024), GPT-4 Achiam et al. (2023), MedLM, Claude 3 Anthropic (2023), OpenBioLLM Pal and Sankarasubbu (2024), Gemini Team et al. (2023), Meditron Chen et al. (2023), and MedLlama Wu et al. (2023a).
@@ 4.2 Quantitative Evaluation
For multiple-choice tasks, accuracy is measured by comparing LLM’s single-letter answer choice [A,B,C,D,E] with the reference. For open-ended questions, semantic similarity is measured using BertScore, comparing the generated response from the language model against a reference answer Zhang* et al. (2020).
@@ 4.3 Qualitative: Human Evaluations and Evaluation Axes
Ratings were on a 5-point scale representing the extent to which the criteria were met. 1 represents “No" or “completely absent", and 5 represents “Yes" or “Absolutely present". Raters were blinded to the answer source (model name or human). Each rater evaluated multiple LLMs or human answers in random blind sequence.
@@ 4.4 Experiment Setup
Checkpoints for open-source models were sourced from HuggingFace and Google’s Vertex AI Studio, while proprietary models were accessed through developers’ API using default hyperparameters. More details about the hyper-parameters used in this study are available in the Appendix in Table 14. Aligning with our community participatory design approach, experiments by multiple collaborators ran on various hardware types including NVIDIA L4, NVIDIA T4, and A100.
@@ Appendix A Appendix
Table 3: Zero-Shot Expert Evaluation of Large Language Models: Accuracy and BERT Scores on Multiple Choice Questions, Short Answer Questions, and Consumer Queries. Some model names are shortened for brevity.
Model MCQ-Expert SAQ Bertscore CQ Bertscore Domain Access Size Training
Gpt-4 0.7568 0.8727 0.8385 General Closed 1.8T Instruct
Gpt-4o 0.7928 0.8825 0.8254 General Closed 12B Instruct
Gpt-3.5 Turbo 0.5629 0.8813 0.0963 General Closed 175B Instruct
Gpt-4o mini 0.7176 0.8808 - General Closed $\sim$8B Instruct
Claude-3.5 Sonnet 0.777 0.8574 - General Closed - Instruct
Claude-3 Sonnet 0.6504 0.8719 0.8141 General Closed 70B Instruct
Claude-3 Opus 0.7455 0.8696 0.8172 General Closed 2T Instruct
Claude 3 Haiku 0.6639 0.8656 0.8163 General Closed 2T Instruct
Gemini Pro 0.6036 0.8601 0.8213 General Closed 540B Instruct
MedLM 0.6036 0.8633 0.8303 Biomedical Closed - Instruct
Gemini Ultra 0.739 0.8716 0.8362 General Closed 1.56T Instruct
MedPalm2 - 0.8716 0.8379 Biomedical Closed 540B Instruct
llama3-405b-instruct-maas 0.7627 - General Open 405B Instruct
Llama3-OpenBioLLM-70B 0.6661 0.8292 0.8283 Biomedical Open 70B Pretrained
Meta Llama3 70B 0.7379 0.7945 0.8372 General Open 70B Instruct
Phi3 Med. 128K 0.6708 0.8661 0.8266 General Open 14B Instruct
Mixtral 8x7B 0.6033 0.8455 0.8259 General Open 46B Instruct
Gemma2 27B 0.6448 0.8617 0.7874 General Open 27B Instruct
Phi3 Mini 128k 0.5903 0.8804 0.8266 General Open 3.8B Pretrained
Phi3 Mini 4k 0.6036 0.874 0.8186 General Open 3.8B Pretrained
Gemma2 9B 0.6153 0.8435 0.8158 General Open 9B Instruct
Llama3 OpenBioLLM 8B 0.4499 0.8629 0.8246 Biomedical Open 8B Pretrained
Llama3 8B 0.4724 0.8804 0.8344 General Open 8B Pretrained
MetaLlama3.1 8B 0.6189 0.8677 - General Open 8B Instruct
PMCLLAMA 7B 0.4629 0.865 - Biomedical Open 7B Finetuned
JSL MedLlama 8B 0.5726 0.8901 0.8303 Biomedical Open 8B Pretrained
Meditron 7B 0.5102 0.8547 - Biomedical Closed 7B Finetuned
BioMistral 7B 0.4402 0.7938 - Biomedical Open 7B Instruct
Mistral 7B v02 0.4847 0.8709 0.8259 General Open 7.2B Instruct
Mistral 7B v03 0.5084 0.8744 - General Open 7.2B Instruct
Gemma2 2B 0.1728 0.8559 0.817 General Open 2B Instruct
@@ E. Onchocerciasis
Meta-Llama-3-8B-Instruct 1 166 0.602 0 11 $0$
Llama3-OpenBioLLM-8B 2 133 1.504 0 5 $0$
Mixtral-8x7B-Instruct-v0.1 1 235 0.426 0 9 $0$
claude-3-haiku 0 83 0.000 0 3 $0$
claude-3-opus 0 79 0.000 0 4 $0$
gemini-ultra 0 36 0.000 0 5 $0$
Llama3-OpenBioLLM-70B 3 169 1.775 0 7 $0$
gemini-pro 0 40 0.000 0 0 $0$
Meta-Llama-3-70B-Instruct 0 150 0.000 0 8 $0$
medlm 1 29 3.448 0 2 $0$
@@
claude-3-opus 1 79 1.266 0 4 $0$
Mixtral-8x7B-Instruct-v0.1 2 235 0.851 0 9 $0$
claude-3-haiku 0 83 0.000 0 3 $0$
gemini-ultra 0 36 0.000 0 5 $0$
medpalm2 0 56 0.000 0 3 $0$
medlm 0 29 0.000 0 2 $0$
@@
gemini-pro 2 40 5.000 0 0 $0$
Meta-Llama-3-70B-Instruct 4 150 2.667 0 8 $0$
Meta-Llama-3-8B-Instruct 3 166 1.807 0 11 $0$
claude-3-opus 1 79 1.266 0 4 $0$
@@
claude-3-haiku 0 83 0.000 0 3 $0$
gemini-ultra 0 36 0.000 0 5 $0$
medlm 0 29 0.000 0 2 $0$

### v4

@@ Abstract
Tobi Olatunji1,2,4 thanks: These authors contributed equally to this work., Charles Nimo2  [^1] Available at:
https://huggingface.co/datasets/intronhealth/afrimedqa_v2
, Abraham Owodunni1,3  [^1] Available at:
https://huggingface.co/datasets/intronhealth/afrimedqa_v2
Tassallah Abdullahi4,12, Emmanuel Ayodele1, Mardhiyah Sanni1,4, Chinemelu Aka1, Folafunmi Omofoye4, Foutse Yuehgoh4, Timothy Faniran4, Bonaventure F. P. Dossou4,10,13, Moshood Yekini4, Jonas Kemp6, Katherine Heller6, Jude Chidubem Omeke4, Chidi Asuzu MD4, Naome A. Etori4,11, Aimérou Ndiaye5, Ifeoma Okoh5, Evans Doe Ocansey5, Wendy Kinara7, Michael Best2, Irfan Essa2,6, Stephen Edward Moore8, Chris Fourie9, Mercy Nyamewaa Asiedu6 thanks: Senior authors. 1Intron, 2Georgia Institute of Technology, 3The Ohio State University, 4BioRAMP, 5Masakhane, 6Google Research, 7Kenyatta University, 8University of Cape Coast, 9SisonkeBiotik, 10MILA Quebec, 11University of Minnesota, 12Brown University 13McGill University tobi@intron.io
@@ 1 Introduction
Feature AfriMed-QA
BioASQ
Krithara et al. (2023)
MedQA-SWE
Hertzberg and Lokrantz (2024)
MedQA
Jin et al. (2020)
MedMCQA
Pal et al. (2022b)
PubMedQA
Jin et al. (2019)
MMLU
Hendrycks et al. (2020)
HealthSearchQA
Singhal et al. (2022a)
@@
To address this gap, we introduce AfriMed-QA, a dataset of 15,275 English, clinically diverse questions and answers, 4,000+ expert multiple-choice questions (MCQs) with answers, over 1,200 open-ended short answer (SAQs) with long-form answers, and 10,000 consumer queries (CQ), to rigorously assess LLM performance for correctness and demographic bias. We evaluate 30 large, small, open, closed, biomedical, and general LLMs using quantitative and qualitative approaches. While the development of the dataset is still in progress, this work establishes a foundation for acquiring diverse and representative health benchmark datasets across LMICs. The dataset is released under a CC-BY-NC-SA 4.0 license.
@@ 3.1 Data Collection
We adapted a web-based platform previously developed by Intron Health 22footnotetext: Intron Health’s biomedical crowd-sourcing platform https://speech.intron.health to crowd-source accented and multilingual clinical speech data at scale across Africa Olatunji et al. (2023). Custom user interfaces were developed to collect each question type, for quality reviews, and for blind human evaluation of LLM responses. The MCQ User Interface and other details about the data collection tool can be found in Appendix Figure 13.
@@ 3.2 Contributor Recruitment and Instructions for Data Collection and Evaluation
Human Evaluations: Consumers provided ratings for LLM responses to CQs on relevance, helpfulness, and local context, but NOT correctness. Due to the higher expertise required, only confirmed clinicians were granted access to projects rating MCQ, SAQ, and CQ answers. Clinician status was confirmed through submitted credentials and background checks reviewing publicly available information about them. Raters were randomly assigned (double-blind) to the answer source (human or LLM). More details on human evaluations are provided in Section 4.3
@@ 4.1 LLM Selection
We evaluate 30 LLMs (Table 3) including open-source and proprietary LLMs, general-purpose and biomedical LLMs, mixture-of-experts, and models of varying sizes from 3B to over 540B parameters. Of these, 13 are proprietary while 17 are open-source. We evaluate LLMs like Phi-3 Abdin et al. (2024), GPT-4 Achiam et al. (2023), MedLM, Claude 3 Anthropic (2023), OpenBioLLM Pal and Sankarasubbu (2024), Gemini Team et al. (2023), Meditron Chen et al. (2023), and MedLlama Wu et al. (2023a).
@@ 4.2 Quantitative Evaluation
For multiple-choice tasks, accuracy is measured by comparing LLM’s single-letter answer choice [A,B,C,D,E] with the reference. For open-ended questions, we evaluate semantic similarity using BERTScore Zhang* et al. (2020) and QuestEval Scialom et al. (2021), both comparing the generated response from the language model against a reference answer, and sentence-level structural overlap using ROUGE-Lsum Lin (2004), which likewise compares the generated response against its reference.
@@ 4.3 Qualitative: Human Evaluations and Evaluation Axes
Ratings were on a 5-point scale representing the extent to which the criteria were met. 1 represents "No" or "completely absent", and 5 represents "Yes" or "Absolutely present". Raters were blinded to the answer source (model name or human). Each rater evaluated multiple LLMs or human answers in random blind sequence.
@@ 4.4 Experiment Setup
Checkpoints for open-source models were sourced from HuggingFace and Google’s Vertex AI Studio, while proprietary models were accessed through developers’ API using default hyperparameters. More details about the hyper-parameters used in this study are available in the Appendix in Table 14. Aligning with our community participatory design approach, experiments by multiple collaborators ran on various hardware types including 1 of either NVIDIA L4, NVIDIA T4, or A100.
@@ 6.2 Localization is still a challenge
6.3 Understanding the domain shift
Although it seems intuitive that medical knowledge is universal, our findings highlight key reasons for regional differences in LLM Performance. Questions require not only general biomedical knowledge but also contextual understanding of African-specific disease patterns, local health challenges, and socio-cultural factors. For example, children in specific communities may be required to take vaccines (e.g. HPV) earlier or later based on logistical access challenges. Diseases may also involve region-specific appearance of symptoms (e.g. appropriate color for skin lesion) and clinical presentation (e.g. cancer patients typically seek professional medical care much later than Western countries, affecting how they are managed on the first visit), and management may reflect available medications, treatment options, and diagnostic equipment. Although there is significant medical knowledge overlap globally, regional variations exist, necessitating physician board certification in each country. Such local variations make questions more challenging for LLMs trained mainly on Western medical data.
@@ 6.7 Performance variation by country
6.8 Limitations of automated metrics for evaluating SAQ and CQ answers
As shown by the narrow range of BERTScore values in Appendix Table 3 (all LLMs cluster between 0.86 and 0.89), its utility in this context is limited—a problem also identified in the literature Hanna and Bojar (2021); Celikyilmaz et al. (2020); Zhang et al. (2024). Appendix Table 3 also reports ROUGE-Lsum and QuestEval scores. ROUGE-Lsum spans a much broader interval (0.009–0.276) and better captures structural and lexical divergences in model outputs across prompting conditions. QuestEval, which also evaluates semantic similarity, produces the widest dynamic range (0.19–0.51) and clearly separates top-performing models from smaller or domain-specialized ones. This variability demonstrates that no single automated metric reliably captures both the correctness and completeness of free-text medical responses Tu et al. (2024), even when metrics are combined. We therefore defer analysis of open-ended answers to future human evaluations where clinicians can better discriminate between semantically similar and clinically correct answers.
@@ 7 Conclusion
8 Ethical Considerations
The public release of healthcare-related datasets often raises important ethical concerns, especially regarding privacy and consent. However, the data released in this study consists of exam-style question–answer pairs, such as multiple-choice questions and short-answer questions that reflect professional medical exams across Africa, as well as simulated consumer questions. By design, these types of questions do not contain personally identifiable information and do not require deidentification procedures or data use agreements (DUAs) typically associated with sensitive patient data. Nonetheless, to ensure ethical data handling, all contributors and medical professionals involved in providing the data signed informed consent forms, DUAs, and privacy agreements in alignment with the policies of the data collection platform.
@@ Appendix A Appendix
Table 3: Zero-Shot Expert Evaluation of Large Language Models: Accuracy, BERT Scores, ROUGE-Lsum (R-L) Scores, and QuestEval (QE) scores, on Multiple Choice Questions, Short Answer Questions, and Consumer Queries. Some model names are shortened for brevity. Additionally, Gen. and Biomed. here is short for General and Biomedical respectively.
Model MCQ Expert SAQ BERT CQ BERT SAQ R-L CQ R-L SAQ QE CQ QE Domain Access Size Type
Gpt-4 0.7568 0.8727 0.8385 0.2054 0.0018 0.4868 0.2352 Gen. Closed 1.8T Instruct
Gpt-4o 0.7928 0.8825 0.8254 0.2519 0.0014 0.4959 0.2352 Gen. Closed 12B Instruct
Gpt-3.5 Turbo 0.5629 0.8813 0.0963 0.2452 0.0015 0.5072 0.2353 Gen. Closed 175B Instruct
Gpt-4o mini 0.7176 0.8808 - 0.2311 - 0.4946 - Gen. Closed $\sim$8B Instruct
Claude-3.5 Sonnet 0.777 0.8574 - - - 0.5071 - Gen. Closed - Instruct
Claude-3 Sonnet 0.6504 0.8719 0.8141 0.1984 0.0010 0.4976 0.2352 Gen. Closed 70B Instruct
Claude-3 Opus 0.7455 0.8696 0.8172 0.1892 0.0010 0.4904 0.2354 Gen. Closed 2T Instruct
Claude 3 Haiku 0.6639 0.8656 0.8163 0.1929 0.0010 0.5004 0.2352 Gen. Closed 2T Instruct
Gemini Pro 0.6036 0.8601 0.8213 0.2061 0.0012 0.4370 0.2347 Gen. Closed 540B Instruct
MedLM 0.6036 0.8633 0.8303 0.1991 0.0014 0.4443 0.2380 Biomed. Closed - Instruct
Gemini Ultra 0.739 0.8716 0.8362 0.2617 0.0018 0.4368 0.2347 Gen. Closed 1.56T Instruct
MedPalm2 - 0.8716 0.8379 0.2253 0.0018 0.4304 0.2352 Biomed. Closed 540B Instruct
llama3-405B 0.7627 - 0.1096 - 0.3744 - Gen. Open 405B Instruct
OpenBioLLM 70B 0.6661 0.8292 0.8283 0.1866 - 0.3107 - Biomed. Open 70B Instruct
Meta Llama3 70B 0.7379 0.7945 0.8372 0.0089 - 0.2331 - Gen. Open 70B Instruct
Phi3 Med. 128K 0.6708 0.8661 0.8266 0.2432 - 0.4999 - Gen. Open 14B Instruct
Mixtral 8x7B 0.6033 0.8455 0.8259 0.2045 - 0.3783 - Gen. Open 46B Instruct
Gemma2 27B 0.6448 0.8617 0.7874 - - - - Gen. Open 27B Instruct
Phi3 Mini 128k 0.5903 0.8804 0.8266 0.2421 0.0014 0.4868 0.2351 Gen. Open 3.8B Pretrained
Phi3 Mini 4k 0.6036 0.874 0.8186 0.2098 0.0011 0.4894 0.2352 Gen. Open 3.8B Pretrained
Gemma2 9B 0.6153 0.8435 0.8158 - - - - Gen. Open 9B Instruct
OpenBioLLM 8B 0.4499 0.8629 0.8246 0.2017 - 0.3877 - Biomed. Open 8B Instruct
Llama3 8B 0.4724 0.8804 0.8344 0.2421 0.0021 0.4868 0.2336 Gen. Open 8B Pretrained
MetaLlama3.1 8B 0.6189 0.8677 - 0.1901 - 0.4817 - Gen. Open 8B Instruct
PMC-Llama 7B 0.4629 0.865 - 0.2194 - 0.3853 - Biomed. Open 7B Finetuned
JSL MedLlama 8B 0.5726 0.8901 0.8303 0.2758 0.0016 0.4478 0.2352 Biomed. Open 8B Pretrained
Meditron 7B 0.5102 0.8547 - 0.1945 - 0.3691 - Biomed. Closed 7B Finetuned
BioMistral 7B 0.4402 0.7938 - 0.2117 - 0.1890 - Biomed. Open 7B Instruct
Mistral 7B v02 0.4847 0.8709 0.8259 0.1989 - 0.4763 0.2355 Gen. Open 7.2B Instruct
Mistral 7B v03 0.5084 0.8744 - 0.2106 - 0.4824 - Gen. Open 7.2B Instruct
Gemma2 2B 0.1728 0.8559 0.817 - - - - Gen. Open 2B Instruct
@@ E. Onchocerciasis
Meta-Llama-3-8B-Instruct 1 166 0.602 0 11 0.000 -
Llama3-OpenBioLLM-8B 2 133 1.504 0 5 0.000 -
Mixtral-8x7B-Instruct-v0.1 1 235 0.426 0 9 0.000 -
claude-3-haiku 0 83 0.000 0 3 0.000 -
claude-3-opus 0 79 0.000 0 4 0.000 -
gemini-ultra 0 36 0.000 0 5 0.000 -
Llama3-OpenBioLLM-70B 3 169 1.775 0 7 0.000 -
gemini-pro 0 40 0.000 0 0 0.000 -
Meta-Llama-3-70B-Instruct 0 150 0.000 0 8 0.000 -
medlm 1 29 3.448 0 2 0.000 -
@@
claude-3-opus 1 79 1.266 0 4 0.000 -
Mixtral-8x7B-Instruct-v0.1 2 235 0.851 0 9 0.000 -
claude-3-haiku 0 83 0.000 0 3 0.000 -
gemini-ultra 0 36 0.000 0 5 0.000 -
medpalm2 0 56 0.000 0 3 0.000 -
medlm 0 29 0.000 0 2 0.000 -
@@
gemini-pro 2 40 5.000 0 0 0.000 -
Meta-Llama-3-70B-Instruct 4 150 2.667 0 8 0.000 -
Meta-Llama-3-8B-Instruct 3 166 1.807 0 11 0.000 -
claude-3-opus 1 79 1.266 0 4 0.000 -
@@
claude-3-haiku 0 83 0.000 0 3 0.000 -
gemini-ultra 0 36 0.000 0 5 0.000 -
medlm 0 29 0.000 0 2 0.000 -